ELECTRONIC SPECTRA OF RhH AND RhD

\underline{C. ZHOU}, W. J. BALFOUR, J. CAO, and C. X. W. QIAN, Department of Chemistry, University of Victoria, Victoria, B.C. Canada V8W 3P6.

Electronic transitions have been observed with LIF techniques in a supersonic molecular beam of He doped with H_{2} or D_{2} and laser-ablated Rh vapor. Bands of RhH have been found near 464.73 and 438.38 nm, which are assigned as (0,0) and (1,0) vibrational bands of a ^{3}Δ_{3} - X^{3}Δ_{3} electronic transition. There is a sudden disapperance of the rotational structure in the (1,0) band of RhH above J′=7 and the spectra associated with higher vibrational levels in the excited states are missing. These observations are attributed to predissociation. Its onset agrees with an ab intio calculation of 22 664-23 955 cm^{−1}^{a} and the estimate of 20 650 ± 1 750 cm^{−1}from thermochemistry.^{b} The RhH(0,0) band also shows irregular Λ-doubling. The corresponding (0,0) and (1,0) transitions in RhD appear at 463.40 and 443.02 nm. Analyses of the spectra indicate that the ground state of RhH is an inverted ^{3}Δ state, which is believed to derive from the electronic configuration ...δ^{3}π^{4}σ^{1}, and provide the following RhH molecular constants(cm^{−1}): B′′_{0}=6.546 [r′′_{0}=0.1610 nm], B′_{0}=6.175, B′_{1}=5.823, ω′′_{e}=1927, ω′′_{e}x′′_{e}=35.7, ω′_{e}=1645, ω′_{e}x′_{e}=174. The observed ground state symmetry and internuclear distance agree with ab initio predictions.^{a} Six additional bands have been found for RhD: near 445.15 nm(Ω′=3 - Ω′′=3), 428.05 nm(Ω′=3 - Ω′′=3), 427.83 nm(Ω′=3 - Ω′′=3), 424.40 nm(Ω′=2 - Ω′′=3), 422.58 nm(Ω′=3 - Ω′′=3), and 409.24 nm(Ω′=4 - Ω′′=3).

aF. Illas, J. Rubio, J. Canellas and J. M. Ricart, J. Chem. Phys., 93, 2603(1990)

bP. B .Armentrout and J. L. Beauchamp, Acc. Chem. Res., 22, 315(1989)